OpenStudy (anonymous):

Solve the equation by factoring. (Do not use mixed numbers in your answers. 3x^2+10x=-3

OpenStudy (anonymous):

you can move the 3 to the other side which gives: 3x^2+10x+3=0 and then do a reverse 'FOIL'. It would end up giving you (3x+1)(x+3)=0 Then find what each x is: 3x=-1 => x=-(1/3) x+3=0 => x=-3

OpenStudy (anonymous):